The distance between Radhas home and her school is 3250 m. Express this distance into km.


 

Grade:9

2 Answers

Dinesh
9 years ago
1000 m = 1km
1 m = 1/1000 km
3250 m = (1/1000) * 3250 km = 3.25 km

Thanks & Regards,
Dinesh Sajnani
AskIITians Faculty

Adarsh
8 years ago
3250m =3250/1000 km (1km = 1000m)
           =3.25km